OpenStudy (anonymous):

Mean, median, and mode of, 58, 58, 75, 86, 94 and 97

OpenStudy (anonymous):

mode=the number showing up the most median= the middle number mean= all the numbers added together then divided by the amount of numbers there.. in this case divided by 6

OpenStudy (amistre64):

id subtract, say 80 from the set and work on the smaller stuff

OpenStudy (amistre64):

or 70 .... 58, 58, 75, 86, 94 and 97 -70 ............................... -12,-12, 5, 16, 24, 27

OpenStudy (amistre64):

then add back 70 to the mean mode etc of these for the original set

OpenStudy (amistre64):

and median will have to be an average of the middle 2 ...... accursed even cardinalities!!

OpenStudy (anonymous):

Thanks Helperock

OpenStudy (anonymous):

You are welcome.

OpenStudy (anonymous):

since there is 6 numbers to find the median you would have to add the 2 middle numbers together then divid them by 2. 75+86=x x/2=median

OpenStudy (anonymous):

so 118 is the median

OpenStudy (amistre64):

the median would be between 75 and 86 ... 118 just doesnt qualify

OpenStudy (anonymous):

no you have to add 75 and 86 together then the answer you get divide that by 2 and you will have the median.

OpenStudy (amistre64):

-12,-12, 5, 16, 24, 27 ^ median (5+16)/2 = 21/2 = 10.5 +70 ; add back the 70 i took out ------ 80.5 to dbl chk: (75+86)/2 = 161/2 = 80.5

OpenStudy (anonymous):

75+86= 161 161/2=80.5 80.5 is your median.